The sum of three times the number and a second number is one. The difference of the first number minus the second number is seven. Find these numbers.
Answer by lwsshak3(11628)

The sum of three times the number and a second number is one. The difference of the first number minus the second number is seven. Find these numbers
..
let x=first number
let y=second number
3x+y=1
x-y=7
y=x-7
3x+x-7=1
4x=8
x=2
y=x-7=2-7=-5
ans:
numbers are: 2 and -5